OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (texaschic101):

y = -2x + 2 In y = mx + b form, the m is the slope. So the slope of this line is -2. A parallel line will have the same slope. y = mx + b m = -2 (4,5) x = 4 and y = 5 now we sub 5 = -2(4) + b 5 = -8 + b 5 + 8 = b 13 = b your parallel equation is : y = -2x + 13